how do u solve 3x +15 when x=12

Found 4 solutions by usyim88hk, gph, solver15, d1234:
Answer by usyim88hk(157)

When x is 12, then 3x+15 can be rewritten as 3*(12)+15
because x is 12, so jsut replace x in 3x+15 by 12 which is now
3(12)+15
= 36+15
= 51

Answer by gph(7)

Subsitute x for 12
Multiply 3*12 = 36
Solve 36 + 15 = 51

Answer by solver15(27)

3x +15 when x=12
plug in the 12 to x
3(12)+15
36+15
=51

Answer by d1234(1)

3x +15 when x=12
3(12)+15
36+15
51
Your answer is 51